2.) x,y (3,1) xy (-3,4)
y=mx+b
3=6x+b=
-b+3=6x
-b=6x-3
b=-6x+3
-1x(-b)=-1x6x-1x(-3)
b=-1x6x-1x(-3)
b=-6-1x(-3)
b=-6x+3
-6/3 = -2
-(6/3)=-2
Rise= 6
Run= 3
I am not sure what I did wrong cam I please get some help??
Find the slope and the y-intercept of the line shown below. Then write its
equation in the slope-intercept form.

Answer by Boreal(15235)

I am assuming the second is (x, y) and not xy.
the slope is -1/2
point slope formula y-y1=m(x-x1) where m slope and (x1, y1) point is
y-1=(-1/2)x (x-3)
y=(-1/2)x+5/2. The slope is -1/2 and the y-intercept is 5/2.
your way 3=6x+b, the x and y you had were reversed, but 6 is not the slope. At this point you don't know the slope.
You can figure it from rise over run, subtracting y1 from y2 or reverse and having that the numerator, then do x1 from x2 or reverse AS LONG AS IT IS IN THE SAME ORDER and get -1/2
now y=mx+b and 1=(-1/2)3+b or 5/2=b, the value above.